OpenStudy (anonymous):

Two coaxial cylindrical conductors are shown in perspective and cross-section above. The inner cylinder has radius a = 2.0 cm, length L = 10.0 m and carries a total charge of Qinner = + 8.0 nC (1 nC = 10-9 C). The outer cylinder has an inner radius b = 6.0 cm, outer radius c = 7.0 cm, length L = 10.0 m and carries a total charge of Qouter = - 16.0 nC (1 nC = 10-9 C). What is Ex, the x-component of the electric field at point P which is located at the midpoint of the length of the cylinders at a distance r = 4.0 cm from the origin and makes an angle of 30.0o with the x-axis?

OpenStudy (anonymous):

picture and wich part of it you don't get?

OpenStudy (anonymous):

https://www.smartphysics.com/Content/Media/Images/EM/04/h4_sphere_v3.png that's the picture and im using the equation \[\frac{ 1 }{ 4\pi \epsilon } \times \frac{ Q inner }{ rL }\] but im not sure if thats the correct equation

OpenStudy (anonymous):

@amistre64